OpenStudy (anonymous):

The cost of printing the school newspaper is $60 for 200 copies and $105 for 400 copies. Write a linear equation modeling this data and find the amount of money it would cost to print 485 copies.

OpenStudy (texaschic101):

200/60 = 3.33 400/105 = 3.80 I am not sure....because if you pay 60 for 200 copies, you are paying 3.33 per copy and if you are paying 105 for 400 copies you are paying 3.80 per copy. I am sorry, I don't know.

OpenStudy (anonymous):

Given two points (60, 200) and (105 , 400) you can make a line. The model of the equation will be Ax + By + C= 0 or you can do y = mx +b

OpenStudy (anonymous):

thanks this helps

OpenStudy (texaschic101):

does he need to find the slope ?

OpenStudy (anonymous):

yes.. it is y2-y1/ x2 - x1.... so its (400-200)/(105-60)..

OpenStudy (texaschic101):

If the slope can reduce, do you reduce it ?

OpenStudy (anonymous):

my answer choices are $124.13 $135 $139.13 $165

OpenStudy (texaschic101):

@jim_thompson5910 ...can you help

jimthompson5910 (jim_thompson5910):

did you guys manage to find the equation?

OpenStudy (anonymous):

Here is what I think the solution is .... m = 4.444.... then plug any of the two points so you choose 400= (4.444)(105) + b so b = -66.66667... To check I will do the other point as well. 200 = (4.444)(60) + b... b= -66.6667 So now all I do is 485 =(4.4444...)(x) -66.666667... solve for x. X = 124.13. So the answer is $124.13. I think....

OpenStudy (anonymous):

im going to try that

jimthompson5910 (jim_thompson5910):

mebs you have the coordinates backwards, which gives you the wrong equation, but you did manage to get the correct answer for the second part

OpenStudy (anonymous):

o swell... excellent my luck always saves me...

OpenStudy (texaschic101):

m = 200/45 = 40/9 y = mx + b ...using points (105,400) 400 = 40/9(105) + b 400 = 4200/9 + b 400 - 4200/9 + b 3600/9 - 4200/9 = b - 600 = b y = 40/9x - 600 I guess I am really way off on this one

jimthompson5910 (jim_thompson5910):

you also have the coordinates backwards I'm afraid

jimthompson5910 (jim_thompson5910):

it should be m = 9/40

OpenStudy (texaschic101):

lol......I am so not good at this

jimthompson5910 (jim_thompson5910):

that's ok, just keep practicing and you'll get better

OpenStudy (texaschic101):

thanks jim....I will

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

so you want the co-ordinates to be (200,60) and (400,105)? What are the correct co-ordinates please explain? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

the coordinates are in the form (x,y) where x = # of copies y = total cost (to make x copies)

jimthompson5910 (jim_thompson5910):

so you have it right

OpenStudy (anonymous):

o goody thanks

jimthompson5910 (jim_thompson5910):

yw